Please help me solve this word problem: a company buys pens at the rate of $6.50 per box for the first 10 boxes, $5.50 per box
for the next 10 boxes, and $4.50 per box for any additional boxes. how many boxes of pens can
be bought for $183.00? Thank you.

Answer by lwsshak3(11628)

Please help me solve this word problem: a company buys pens at the rate of $6.50 per box for the first 10 boxes, $5.50 per box
for the next 10 boxes, and $4.50 per box for any additional boxes. how many boxes of pens can
be bought for $183.00?
***
let x=number of boxes of pens which can be bought for $183.
10(6.50)+10(5.50)+(x-20)(4.50)=$183.00
65+55+4.5x-90)=183
120+4.5x-90=183
4.5x=153
x=153/4.5=34
number of boxes of pens which can be bought for $183=34